# Manifolds, Random Matrices and Spectral Gaps: The geometric phases of generative diffusion

### v1

@@ 5.3 Phase III: Manifold consolidation
where $\Pi=F(F^{\top}F)^{-1}F^{\top}$ is the projection matrix over the manifold.
Figure 2: Visualization of the trivial (I), coverage (II) and consolidation phase (III).
The component of the score orthogonal to the manifold diverges proportionally to $t^{-1}$, while the tangent components converge to a constant and become therefore negligible in this regime. This results in the consolidation of the gap corresponding to the manifold dimensionality $m$ and to the (relative) closure of the intermediate gaps. Therefore, in this final phase the dynamics of the model simply projects the particles into the manifold $\mathcal{M}_{t}\rightarrow\mathcal{M}_{0}$. In the generative modeling literature, this phenomenon is also known as manifold overfitting as the terms corresponding the the internal distribution is negligible (Loaiza-Ganem et al.,, 2022). Fortunately, this internal coverage occurs in the model throughout intermediate phases across the previous diffusion process, resulting in a balanced coverage of the internal distribution.
6 Theoretical analysis of the spectral of linear diffusion models
@@ 6.1 Linear manifolds
Normally the distribution $p_{0}(\boldsymbol{x})$ is unknown. It is however interesting to investigate tractable special case where the distribution is a multivariate Gaussian defined on a linear manifold:
$\boldsymbol{y}^{\mu}=F\boldsymbol{z}^{\mu}$ (10)
where $F\in\mathbb{R}^{d\times m}$ is an arbitrary projection matrix that implicitly define the structure of the latent manifold. and $\boldsymbol{z}^{\mu}\sim\mathcal{N}(0,I_{m})$ the latent space vector. In this setting, the distribution can be explicitly written as
$p_{0}(\mathbf{x})=\mathbb{E}_{\mathbf{z}\sim\mathcal{N}(0,I_{m})}\delta(\mathbf{x}-F\mathbf{z})=\mathcal{N}(\mathbf{x};0,FF^{\top})$ (11)
Therefore, the density of the process at a given time $t$ is again Gaussian and can be computed from
@@
While linear manifolds are very simple when compared with real data, they still exhibit a rich and non-trivial phenomenology that elucidate several universal phenomena of diffusion under the manifold hypothesis. In fact, these linear models capture the structure of tangent spaces of smooth manifolds.
@@ 6.2 The linear random-manifold score in large dimensions: A random matrix analysis
In the following, we provide an outline of our theoretical results on the distribution of spectral gaps in the matrix $J_{t}$ under random linear manifolds. This choice reflects the fact that the distribution and support of the data is usually not known in advance, and it is therefore important to quantify the statistical variability induced by this uncertainty. Specifically, we will consider random projection matrices $F$ distributed according to different distributions. To ensure tractability, we perform the analysis in the limit of large $d$ (visible) and $m$ (latent) dimensions while keeping the ratio $\alpha_{m}=m/d$ constant.
@@ 6.3 The isotropic case
If the elements of the projection matrix $F$ are sampled as $F_{ij}\sim\mathcal{N}(0,\sigma^{2}/m)$, we are able to derive analytically the full expression of the distribution of the eigenvalues of $J_{t}$. It is given by a simple transformation of the distribution of the eigenvalues of $F^{\top}F$, which is known to be the Marchenko-Pastur distribution reported in Appendix A.1.
(a) t = 10
(b) t = 1
(c) t = 0.01
(d) t = 10
(e) t = 1
(f) t = 0.01
@@ 6.4 Intermediate gaps and subspaces with different variances
Another relevant case for our study is the one where we consider a manifold having multiple subspaces with different variances. Here we will focus on the instance of two distinct variances. This scenario is reproduced by considering a number $f\cdot m$ of columns of $F$ to have element Gaussian distributed with zero mean and variance $\sigma_{1}^{2}/m$, and the remaining $(1-f)\cdot m$ columns with elements from a Gaussian with variance $\sigma_{2}^{2}/m$. The spectrum of $J_{t}$ can be computed also in this case, as explained in Appendix A.2. The density function of the eigenvalues shows a transient behavior of the spectrum in the form of an intermediate drop in the estimated dimensionality of the hidden data manifold. This behavior is reported in Fig. 4. Even though the expression of the spectral density doesn’t have an explicit analytical form and has to be computed numerically, one can adopt a special assumption on the behavior of the density of the eigenvalues of $F^{\top}F$ to estimate the typical times at which the intermediate drop occurs. Generally speaking, the spectrum of $F^{\top}F$ can be composed of two separated bulks, as observable in Fig. (9). This happens when $\sigma_{2}^{2}$ and $\sigma_{1}^{2}$ are significantly different. In analogy with the single variance scenario, we name $\gamma_{+}$ the left bound of the bulk associated to higher eigenvalues, i.e. with the higher variance, and $\gamma_{-}$ the right bound of the bulk associated with smaller eigenvalues, i.e. smaller variance. Most commonly, $\gamma_{-}=\gamma_{-}(\sigma_{2})$ and $\gamma_{+}=\gamma_{+}(\sigma_{1})$. In this case the gap-forming phenomenology provides for two distinct processes: the emergence of intermediate gaps (i.e. steps in the dimensionality plot) between spearated bulks of the spectrum, the opening of a final gap that allows to infer the dimensionality of the full manifold. The width of the intermediate gap between two bulks can be obtained from Eq. (27) as
@@
$t_{\text{in}}(\Delta)$ $\simeq\Delta^{-1}f\left(1-\sqrt{\frac{1}{f\alpha_{m}}}\right)^{2}\sigma_{1}^{2},$ (24)
$t_{\text{fin}}(\Delta)$ $\simeq\Delta(1-f)\left(1+\sqrt{\frac{1}{(1-f)\alpha_{m}}}\right)^{2}\sigma_{2}^{2},$ (25)
@@
This same analysis can be extended to the more general case where the spectral density is known to be formed by different detached bulks, associated to hierarchically smaller variances of the data. The evolution of the intermediate gaps in a double-variance diffusion model is reported in Fig. 4: notice that $t_{\text{max}}=\mathcal{O}(\sigma_{2})$ is consistent with Fig. 4b and 4f, where the gap was found to be maximum in width. It is worth noting that subspaces with higher variances are the first ones to be explored by diffusion, and to be learned by the model. This point suggests that the model is sensitive to the parameters of the probability distribution on the manifold as recently suggested by other works in the literature (see section 9 for further details).
@@ 7 Experiments with synthetic linear datasets
We first measure the spectrum of the singular values of the Jacobian of a score function trained through a neural network on a linear manifold data-model generated by two variances $\sigma_{1}^{2}$, $\sigma_{2}^{2}$ as described in Section 6.4. Results are reported in Figure 5 (left and central panels). The opening of the gaps is consistent with the theory for the exact score: an intermediate gap associated to the subspace with higher variance first opens; subsequently, the gap relative to the lower variance subspace, which here corresponds to the final gap, opens. We can infer the dimensions of the subspaces by subtracting the location of the dashed vertical lines to $d$. We underline the fact that higher-variance subspaces are learned first by repeating the experiment after swapping the values of the variances. Eventually, the right panel in Figure 5 reports the same experiments where variances are uniformly generated in the interval $[10^{-3},1]$: it is evident that the $d$ intermediate gaps is now a continuous line, as it is expected to be in more realistic natural data-sets.
@@
Figure 5: Ordered singular values obtained with the trained score model, for difference variances on the subspaces. Data are generated according to the linear-manifold model with $d=100$ and $m=40$. Left: $\sigma_{1}^{2}=1$, $\sigma_{2}^{2}=0.01$, $f=0.75$; Center: $\sigma_{1}^{2}=0.01$, $\sigma_{2}^{2}=1$, $f=0.75$; Right: variances sampled uniformly between $10^{-3}$ and $1$.
@@
Figure 6: Comparison between spectra obtained with the trained score model and with the numerical analysis, for difference variances on the subspaces. Data are generated according to the linear-manifold model with $d=100$ and $m=40$, $\sigma_{1}^{2}=1$, $\sigma_{2}^{2}=0.01$, $f=0.75$; from left to right, the spectra are evaluated a time $t\approx 0.45$, $t\approx 0.3$, $t\approx 0.11$.
@@
We will now compare the gaps computed analytically with ones obtained from real neural networks trained on the same linear data-model. The results of such comparison are presented in Figure 6, and they show a good agreement between the ordered distribution of the singular values obtained through empirical methods, and the relative analytical counterpart, computed through the replica method. The opening of the predicted intermediate gaps signal the right dimension of the linear subspaces as verified from the experiments. One can notice from the figure that the analytical profile shows the shape of a sharp step between the zero value along the $x$-axis and the first appearing gap: this shape is related to the Dirac-delta spike that the spectrum of the eigenvalues presents at $r=-1$ (see Appendix A.2 for details about the spectrum); on the other hand, the numerical profile looks different in the same region, and this behaviour is associated to the absence of the spike in the distribution of the singular values, that leaves room to a separated bulk from the other ones. This evident discrepancy between theory and experiment is probably due to the final configuration of the trained neural network and leaves space for further investigations.
@@ 8 Experiments with natural image datasets
While our theoretical analysis is limited to linear random-manifold models, several qualitative figures of its phenomenology can be observed in networks trained in natural images. Fig. 7 shows the temputal evolution of the spectrum estimated numerically from the Jacobian of models trained on MNIST, Cifar10 and CelebA and averaged over $100$ images. The details of this experiment are given in Supp. C and Supp. B. We see that for low times (at the end of the reverse process), the network trained on MNIST shows a sharp total manifold gap at around the $700$-th singular value. Consistently with our theory, for larger time the spectrum becomes smoother. The non averaged result obtained around single images also reveal the presence of several discernible small intermediate gaps (see left panel in Fig. 7). The results on Cifar10 and CelebA are more difficult to interpret, probably due to a less well-defined manifold and subspace structure. However, the inflection points of the spectra reveal a more suble gap structure, which recedes for small values of $t$ as more local subspaces are revealed.
Figure 7: Jacobian spectra of diffusion models trained on MNIST, Cifar10 and CelebA.
@@ 9 Related work & Discussion
The evolution of the fixed-points of the exact score was studied in (Raya and Ambrogioni,, 2024) for the analysis spontaneous symmetry breaking and in (Biroli and Mézard,, 2023) for the analysis of memorization and glassy phase transitions. The use of spectral gaps to quantify the dimensionality of the manifold was introduced in (Stanczuk et al.,, 2022), where the total manifold gap is analyzed. Several recent studies investigated the local linear structure of trained diffusion models. For example, (Kadkhodaie et al.,, 2024) studied the expansion of the Jacobian of trained models and described it as an optimal geometry-adaptive Harmonic representation. Similarly, (Chen et al., 2024b, ) characterized the linear expansion of the Jacobian of trained networks and characterized the resulting components in terms of their frequency content. Our work can be seen as a theoretical complement to this more applied line of research, as we provide a comprehensive random-matrix analysis of the phenomenon in tractable models. The dynamic geometry of diffusion manifolds were also investigated in (Chen et al., 2024a, ) using techniques inspired by research on latent generators such as GANs and autoencoders. Another recent work (W. et al.,, 2024) uses parameterized low-rank score functions to show an equivalence between training diffusion model and subspace clustering. In our work, we prove results for the exact score for high-dimensional datasets. Finally, (Sakamoto et al.,, 2024) studies the dynamic geometry of tubular neighborhoods of the latent manifold and connect their singularities with spontaneous symmetry breaking events. Generative diffusion models exhibit rich geometric structures that have the potential to explain their impressive generative capabilities. Our work introduces the use of random-matrix theory techniques for the analysis of their dynamic local geometry and gives room to the use of advanced statistical physics techniques, which may in the future unveil more global, topological and non-linear aspects of the dynamic geometry of diffusion generative models.

### v4

@@ 5.3 Phase III: Manifold consolidation
where $\Pi=F(F^{\top}F)^{-1}F^{\top}$ is the projection matrix over the manifold. The component of the score orthogonal to the manifold diverges proportionally to $t^{-1}$, while the tangent components converge to a constant and become therefore negligible in this regime. This results in the consolidation of the gap corresponding to the manifold dimensionality $m$ and to the (relative) closure of the intermediate gaps. Therefore, in this final phase the dynamics of the model simply projects the particles into the manifold $\mathcal{M}_{t}\rightarrow\mathcal{M}$. In the generative modeling literature, this phenomenon is also known as manifold overfitting as the terms corresponding the internal distribution is negligible (Loaiza-Ganem et al.,, 2022). In the next section we comment on how these three phases can explain why diffusion models are not affected by this phenomenon, namely why we claim that the manifold is consolidated rather than overfitted.
5.4 The geometric phases and manifold overfitting
The probability density of data defined on a manifold is a spiked object $\delta_{\mathcal{M}}(x)\rho(x)$, where the Dirac-delta $\delta_{\mathcal{M}}(x)$ determines the manifold and $\rho(x)$ determines its internal density. Likelihood-based generative models are defined by a highly parameterized likelihood function $f(\boldsymbol{x};\boldsymbol{\theta})$, whose parameters are trained by minimizing the loss
which maximizes the probability of the data given the model. This maximum likelihood loss is minimized if $f(\boldsymbol{x};\boldsymbol{\theta})=p_{0}(\boldsymbol{x})$. A trained likelihood-based model can only fit the true density by having it to diverge to infinity on the manifold. Such a divergence makes it impossible to correctly model the internal density $\rho(x)$. More problematically, the optimization problem becomes almost insensitive to the internal density $\rho(\boldsymbol{x})$. This phenomenon is called manifold overfitting (Loaiza-Ganem et al.,, 2022), since the trained model fits the manifold while ignoring its internal density, resulting in poor generation.
Our analysis suggests that the temporal dynamics of generative diffusion models overcome this limitation because, for intermediate values of $t$, the score is still sensitive to the density internal to the manifold, which can be identified through the differences in the tangent singular values. During this manifold coverage phase, the score directs the dispersion of the particles according to these differences, with higher singular values resulting in larger ‘opposing force’ from the score, which results in smaller displacements of the generated samples along these directions. For $t$ tending to zero, these differences are suppressed due to the divergence of the likelihood, which results in a score function that is orthogonal to the manifold and that is insensitive to $\rho(x)$. However, at this stage of generative diffusion the internal dispersion of the particles have already been affected by the previous coverage phase and therefore the manifold overfitting of the score does not negatively affect generation. Instead, the consolidation phase plays the important role of projecting the particles to the support of the data.
6 Theoretical analysis of the spectral gaps in linear diffusion models
@@ 6.1 Linear manifolds
Normally the distribution $p_{0}(\boldsymbol{x})$ is unknown. It is however interesting to investigate tractable special case where the distribution is a multivariate Gaussian defined as in Eq. 8 where $F\in\mathbb{R}^{d\times m}$ is an arbitrary projection matrix that implicitly define the structure of the latent manifold. and $\boldsymbol{z}^{\mu}\sim\mathcal{N}(0,I_{m})$ the latent space vector. In this setting, the distribution can be explicitly written as $p_{0}(\mathbf{x})=\mathbb{E}_{\mathbf{z}\sim\mathcal{N}(0,I_{m})}\delta(\mathbf{x}-F\mathbf{z})=\mathcal{N}(\mathbf{x};0,FF^{\top})$. Therefore, the density of the process at a given time $t$ is again Gaussian and can be computed from
@@
While linear manifolds are very simple when compared with real data, they still exhibit a rich and non-trivial phenomenology that elucidate several universal phenomena of diffusion under the manifold hypothesis. In fact, these linear models capture the structure of tangent spaces of smooth manifolds (see Supp. D).
@@ 6.2 The linear random-manifold score in large dimensions: A random matrix analysis
In the following, we provide an outline of our theoretical results on the distribution of spectral gaps in the matrix $J_{t}$ under random linear manifolds. This choice reflects the fact that the distribution and support of the data are usually not known in advance, and it is therefore important to quantify the statistical variability induced by this uncertainty. Specifically, we will consider random projection matrices $F$ distributed according to different distributions. To ensure tractability, we perform the analysis in the limit of large $d$ (visible) and $m$ (latent) dimensions while keeping the ratio $\alpha_{m}=m/d$ constant.
@@ 6.3 The isotropic case
If the elements of the projection matrix $F$ are sampled as $F_{ij}\sim\mathcal{N}(0,\sigma^{2}/m)$, we are able to derive analytically the full expression of the distribution of the eigenvalues of $J_{t}$. It is given by a simple transformation of the distribution of the eigenvalues of $F^{\top}F$, which is known to be the Marchenko-Pastur distribution reported in Supp. A.1.
@@ 6.4 Intermediate gaps and subspaces with different variances
Another relevant case for our study is the one where we consider a manifold having multiple subspaces with different variances. Here we will focus on the instance of two distinct variances. This scenario is reproduced by considering a number $f\cdot m$ of columns of $F$ to have element Gaussian distributed with zero mean and variance $\sigma_{1}^{2}/m$, and the remaining $(1-f)\cdot m$ columns with elements from a Gaussian with variance $\sigma_{2}^{2}/m$. The spectrum of $J_{t}$ can be computed also in this case, as explained in Supp. A.2. The density function of the eigenvalues shows a transient behavior of the spectrum in the form of an intermediate drop in the estimated dimensionality of the hidden data manifold. This behavior is reported in Fig. 2. Even though the expression of the spectral density doesn’t have an explicit analytical form and has to be computed numerically, one can adopt a special assumption on the behavior of the density of the eigenvalues of $F^{\top}F$ to estimate the typical times at which the intermediate drop occurs. Generally speaking, the spectrum of $F^{\top}F$ can be composed of two separated bulks, as observable in Fig. 8. This happens when $\sigma_{2}^{2}$ and $\sigma_{1}^{2}$ are significantly different. In analogy with the single variance scenario, we name $\gamma_{+}$ the left bound of the bulk associated with higher eigenvalues, i.e. with the higher variance, and $\gamma_{-}$ the right bound of the bulk associated with smaller eigenvalues, i.e. smaller variance. Most commonly, $\gamma_{-}=\gamma_{-}(\sigma_{2})$ and $\gamma_{+}=\gamma_{+}(\sigma_{1})$. In this case the gap-forming phenomenology provides for two distinct processes: the emergence of intermediate gaps (i.e. steps in the dimensionality plot) between separated bulks of the spectrum, the opening of a final gap that allows to infer the dimensionality of the full manifold. The width of the intermediate gap between two bulks can be obtained from Eq. (26) as
@@
$t_{\text{in}}(\Delta)\simeq\Delta^{-1}f\left(1-\sqrt{\frac{1}{f\alpha_{m}}}\right)^{2}\sigma_{1}^{2},\qquad t_{\text{fin}}(\Delta)\simeq\Delta(1-f)\left(1+\sqrt{\frac{1}{(1-f)\alpha_{m}}}\right)^{2}\sigma_{2}^{2},$ (24)
@@
This same analysis can be extended to the more general case where the spectral density is known to be formed by different detached bulks, associated with hierarchically smaller variances of the data. The evolution of the intermediate gaps in a double-variance diffusion model is reported in Fig. 2: notice that $t_{\text{max}}=\mathcal{O}(\sigma_{2})$ is consistent with Fig. 2b and 2f, where the gap was found to be maximum in width. It is worth noting that subspaces with higher variances are the first ones to be explored by diffusion, and to be learned by the model. This point suggests that the model is sensitive to the parameters of the probability distribution on the manifold as recently suggested by other works in the literature (see Section 9 for further details).
@@ 7 Experiments with synthetic linear datasets
We first measure the spectrum of the singular values of the Jacobian of a score function trained through a neural network on a linear manifold data-model generated by two variances $\sigma_{1}^{2}$, $\sigma_{2}^{2}$ as described in Section 6.4. Results are reported in Fig. 3 (left and central panels). The opening of the gaps is consistent with the theory for the exact score: an intermediate gap associated with the subspace with higher variance first opens; subsequently, the gap relative to the lower variance subspace, which here corresponds to the final gap, opens. We can infer the dimensions of the subspaces by subtracting the location of the dashed vertical lines from $d$. We underline the fact that higher-variance subspaces are learned first by repeating the experiment after swapping the values of the variances. Eventually, the right panel in Fig. 3 reports the same experiments where variances are uniformly generated in the interval $[10^{-2},1]$: it is evident that the $d$ intermediate gaps is now a continuous line, as it is expected to be in more realistic natural data-sets.
@@
Figure 3: Ordered singular values obtained with the trained score model, for different variances on the subspaces. Data are generated according to the linear-manifold model with $d=100$ and $m=40$. Left: $\sigma_{1}^{2}=1$, $\sigma_{2}^{2}=0.01$, $f=0.75$; Center: $\sigma_{1}^{2}=0.01$, $\sigma_{2}^{2}=1$, $f=0.75$; Right: a progressive number $n$ of variances sampled uniformly between $10^{-2}$ and $1$ each one assigned to a fraction $f=1/n$ of matrix columns. The neural network is trained as prescribed in Supp. B and spectra are measured according to Supp. C.
@@
Figure 4: Comparison between spectra obtained with the trained score model and with the numerical analysis, for different variances on the subspaces. Data are generated according to the linear-manifold model with $d=100$ and $m=40$, $\sigma_{1}^{2}=1$, $\sigma_{2}^{2}=0.01$, $f=0.75$; from left to right, the spectra are evaluated a time $t\approx 0.45$, $t\approx 0.3$, $t\approx 0.11$. The neural network is trained as prescribed in Supp. B and spectra are measured according to Supp. C.
@@
We will now compare the gaps computed analytically with ones obtained from real neural networks trained on the same linear data-model. The results of such comparison are presented in Fig. 4, and they show a good agreement between the ordered distribution of the singular values obtained through empirical methods, and the relative analytical counterpart, computed through the replica method. The opening of the predicted intermediate gaps signal the right dimension of the linear subspaces as verified from the experiments. One can notice from the figure that the analytical profile shows the shape of a sharp step between the zero value along the $x$-axis and the first appearing gap: this shape is related to the Dirac-delta spike that the spectrum of the eigenvalues presents at $-1$ (see Supp. A.2 for details about the spectrum); on the other hand, the numerical profile looks different in the same region, and this behavior is associated with the absence of the spike in the distribution of the singular values, that leaves room to a separated bulk from the other ones. This evident discrepancy between theory and experiment is probably due to the final configuration of the trained neural network and leaves space for further investigations.
@@ 8 Experiments with natural image datasets
While our theoretical analysis is limited to linear random manifold models, several qualitative aspects of its phenomenology can be observed in networks trained on natural images. Fig. 5 shows the temporal evolution of the spectrum estimated numerically from the Jacobian of models trained on MNIST, Cifar10 and CelebA. Details about the training process are provided in Supp. B and Supp. C. In these experiments we can recognize the three geometric phases of diffusion described above:
  1. 1.
    Trivial phase: at large times (i.e. from $t=200$ to $t=100$) the ordered spectrum of the singular values appears flat, suggesting the diffusive motion to be Brownian in the ambient space.
  2. 2.
    Manifold coverage phase: at intermediate times the spectrum shows a clear trace of multiple simultaneous opening gaps. The shape of the curves is, however, different from the controlled scenario showed in Section 6, due to two different reasons: similar latent variances associated to different latent dimensions imply a smoothing of the curve, as displayed by Fig. 3 (right panel); the local eigenspace of the data is complex and hard to model microscopically: for instance, the pixellated appearance of Cifar10 images might explain the scarce emergence of the gaps, while the larger gap structure showed by CelebA might be due to correlations among the latent variances.
  3. 3.
    Manifold consolidation phase: at small times (below $t=2$) we finally see only the full manifold gap open and progressively sharpening, in the sense that many singular values become exactly zero. The spectra of this phase are the only ones analyzed in Stanczuk et al., (2022) and used to estimate the manifold dimensionality. For instance, we see from Fig. 5 (left panel), that at the end of the reverse process the network trained on MNIST shows a latent dimension $m\sim 100$ that is coherent with Stanczuk et al., (2022).
We conclude that, although our analysis focuses on the local structure of the data manifold (or the stable latent set in diffusion time), it is effectively supported by experiments on real-world complex datasets.
Figure 5: Jacobian spectra of diffusion models trained on MNIST, Cifar10 and CelebA. The neural network is trained as prescribed in Supp. B and spectra are measured according to Supp. C.
@@ 9 Related work & Discussion
The evolution of the fixed-points of the exact score was studied in (Raya and Ambrogioni,, 2024) for the analysis of the spontaneous symmetry phenomenon breaking and in (Biroli and Mézard,, 2023) for the analysis of memorization and glassy phase transitions. The use of spectral gaps to quantify the dimensionality of the manifold was introduced in (Stanczuk et al.,, 2022), where the total manifold gap is analyzed. Several recent studies investigated the local linear structure of trained diffusion models. For example, (Kadkhodaie et al.,, 2024) studied the expansion of the Jacobian of trained models and described it as an optimal geometry-adaptive Harmonic representation. Similarly, (Chen et al., 2024b, ) characterized the linear expansion of the Jacobian of trained networks and characterized the resulting components in terms of their frequency content. Our work can be seen as a theoretical complement to this more applied line of research, as we provide a comprehensive random-matrix analysis of the phenomenon in tractable models. The dynamic geometry of diffusion manifolds was also investigated in (Chen et al., 2024a, ) using techniques inspired by research on latent generators such as GANs and autoencoders. Another recent work (Wang et al.,, 2024) uses parameterized low-rank score functions to show an equivalence between training diffusion model and subspace clustering. In our work, we prove results for the exact score for high-dimensional datasets. Finally, (Sakamoto et al.,, 2024) studies the dynamic geometry of tubular neighborhoods of the latent manifold and connect their singularities with spontaneous symmetry breaking events. Generative diffusion models exhibit rich geometric structures that have the potential to explain their impressive generative capabilities. Our work introduces the use of random-matrix theory techniques for the analysis of their dynamic local geometry and paves the way for the use of advanced statistical physics techniques, which may in the future unveil more global, topological and non-linear aspects of the dynamic geometry of diffusion generative models.